# VecCity: A Taxonomy-guided Library for Map Entity Representation Learning [Experiment, Analysis & Benchmark]

### v1

@@ 2.2.2. Graph-based Models
Methods Encoders Pretraining Tasks Map Entities Features Aux. Data
Token Graph Seq TokRI TRCL AToCL NFI GAu NCL AGCL TrajP MTR ATrCL POI RS LP Sem Spa Temp Traj Rel
Point of Interest CWAP (Liu et al., 2016) CIT
CAPE (Chang et al., 2018) CIT
SkipG* (Mikolov et al., 2013) CIT
Tale* (Wan et al., 2021) CIT
Teaser* (Zhao et al., 2017) CIT GR
Hier* (Shimizu et al., 2020) CIT GR
P2Vec* (Feng et al., 2017) CIT GR
CACSR* (Gong et al., 2023) CIT
CTLE* (Lin et al., 2021) CIT
Road Segment TrajRNE (Schestakov et al., 2023) GR/SR
DyToast (Chen et al., 2024b) CDT GR
JGRM (Ma et al., 2024) CDT GR
RN2Vec* (Wang et al., 2020) GR
HRNR* (Wu et al., 2020) GR
SARN* (Chang et al., 2023) GR
Toast* (Chen et al., 2021) CDT GR
HRoad* (Zhang and Long, 2023) GR
START* (Jiang et al., 2023) CDT GR
JCLRNT* (Mao et al., 2022) CDT GR
Land Parcel CGAL (Zhang et al., 2019a) GR/SR
DLCL (Du et al., 2019) GR/SR
HUGAT (Kim and Yoon, 2022) GR/SR
Re2Vec (Luo et al., 2022) GR/SR
ReDCL (Li et al., 2023) GR
HAFus (Sun et al., 2024) GR/SR
ReCP (Li et al., 2024a) SR
ZEMob* (Yao et al., 2018) SR
GMEL* (Liu et al., 2020) GR/SR
MGFN* (Wu et al., 2022) SR
HDGE* (Wang and Li, 2017) GR/SR
MVURE* (Zhang et al., 2020) SR
ReMVC* (Zhang et al., 2022) SR
HREP* (Zhou et al., 2023) GR/SR
@@ 2.2.3. Sequence-based Models
Encoders. Sequence-based encoders aim to capture temporal dependencies among map entities within trajectory-type auxiliary data. They use sequential deep learning models, such as LSTMs (Sak et al., 2014), GRUs (Cho, 2014), or Transformers (Vaswani et al., 2017), as backbones to encode the embeddings of map entities along the same trajectory into a sequence of representation vectors. These input embeddings are typically generated by token-based or graph-based encoders. Specifically, for map entities along a trajectory $tr$, let their embedding sequence be $\bm{R}_{tr}=(\bm{r}_{tr_{1}},\ldots,\bm{r}_{tr_{k}},\ldots,\bm{r}_{tr_{K}})$. The sequence-based model generates representation vectors for these entities as:
(3) $\small\bm{S}_{tr}=(\bm{s}_{tr_{1}},\ldots,\bm{s}_{tr_{k}},\ldots,\bm{s}_{tr_{K}})=\mathrm{SEQ}(\bm{R}_{tr}),$
where $\mathrm{SEQ}(\cdot)$ is the sequence-based encoder, and $\bm{s}_{tr_{k}}$ is the representation vector for the $k$-th entity along the trajectory $tr$.
In trajectory data, each location sample is typically associated with a timestamp. Integrating temporal information from timestamps into entity representations has become a key focus of sequence-based encoders. Early methods divide time into discrete slots, such as the hour of the day or the day of the week, and use these slot indices to generate input embeddings using token-based encoders (Mao et al., 2022; Feng et al., 2017; Zhao et al., 2017). More advanced models embed timestamps within the components of sequential models. For example, JCLRNT (Mao et al., 2022), Toast (Chen et al., 2021), and CTLE (Lin et al., 2021) incorporate timestamps into the positional encoding of Transformers. Similarly, START (Jiang et al., 2023) integrates the travel time of trajectory through a road segment into the attention of Transformers.
@@
  • Trajectory Prediction (TrajP). The goal of this task is to predict the second half of a trajectory using the first half. Given a trajectory $tr$, this task is formally defined as $\hat{tr}_{[k+1:K]}=Pre(\bm{S}_{[1:k]}),~{}\bm{S}_{[1:k]}=Enc(tr_{[1:k]})$, where $Enc(\cdot)$ is the sequence-based encoder, and $Pre(\cdot)$ is the prediction function. Here, $tr_{[1:k]}$ represents the first $k$ samples of the trajectory, and $\bm{S}_{[1:k]}$ is their corresponding sequence of representation vectors. The objective is to minimize the distance between $\hat{tr}_{[k+1:K]}$ and the ground truth.
  • Masked Trajectory Recovery (MTR). Inspired by the masked language model (MLM) used in pre-trained language models (Devlin et al., 2019; Liu et al., 2019), this task randomly masks part of the samples in a trajectory and employs an autoencoder to recover them. Given a trajectory $tr$ and its masked version $\tilde{tr}$, the recovery process is defined as $\hat{tr}=Dec(\tilde{\bm{S}}),~{}\tilde{\bm{S}}=Enc(\tilde{tr})$, where $Enc(\cdot)$ is the sequence-based encoder, and $Dec(\cdot)$ is the decoder, implemented using a deep model. The recovered trajectory is denoted as $\hat{tr}$. The objective is to minimize the distance between $tr$ and $\hat{tr}$.
  • Augmented Trajectory Contrastive Learning (ATrCL). This task creates multiple augmented versions of an original trajectory using data augmentation techniques (Fang et al., 2020; Zhao et al., 2019). Sequence-based encoders are then used to generate representations for both the original and augmented trajectories. A contrastive learning loss is applied to increase the similarity between representations derived from the same original trajectory while decreasing the similarity between those from different trajectories.
In TrajP tasks, the prediction labels can be various. For example, Hier (Shimizu et al., 2020) uses entity IDs as prediction labels, while CTLE (Lin et al., 2021) employs visiting time as labels to incorporate temporal information. In MTR tasks, the main difference across methods lies in how masking positions are chosen. A basic approach is to randomly mask positions within trajectories (Lin et al., 2021), but this overlooks the sequential dependencies between samples. To address this, START (Jiang et al., 2023) and Toast (Chen et al., 2021) mask continuous sub-sequences within a trajectory. In ATrCL tasks, different methods adopt distinct augmentation techniques to enhance trajectory representations. JCLRNT (Mao et al., 2022) generates augmented trajectories by randomly deleting or replacing samples within the trajectory. Considering contextual information, START (Jiang et al., 2023) masks a continuous sub-sequence or replaces it with a new one. CACSR (Gong et al., 2023) introduces white noise into the representations to augment trajectories.
@@ 2.2.4. Module Integration of Encoders and Pre-training Tasks
The encoders and pre-training tasks introduced in the previous sections form the fundamental modules for MapRL models. A complete MapRL model can integrate multiple types of encoders and pre-training tasks, leveraging their complementary strengths to enhance representation learning. For the encoder modules, existing models typically follow a “token $\rightarrow$ graph $\rightarrow$ sequence” modeling pipeline. This pipeline first extracts basic features using token-based encoders, then captures relational structures through graph-based encoders, and finally models temporal dependencies with sequence-based encoders. For encoders at each stage of the pipeline, the MapRL model can simultaneously employ multiple encoders for different pre-training tasks. Table 1 summarizes the combinations of encoders and pre-training tasks used in mainstream models.
For the three-stage encoder pipeline, MapRL models follow two pre-training paradigms. $i$) Sequential Training: The encoders are trained in stages – starting with token-based, followed by graph-based, and finally sequence-based – each refining the previous stage’s output. $ii$) Joint Training: All encoders are optimized together using multi-task learning, capturing features, relations, and temporal patterns in an integrated manner. These paradigms offer flexibility in model impanelments, depending on the specific requirements of the MapRL task.
@@ 2.3. Downstream Tasks
Tasks for POIs. Common downstream tasks for POI entities fall into three categories:
  • POI Classification (POIC): This task predicts the category of a POI based on its representation. The downstream model typically uses an MLP network with a SoftMax classifier.
  • Next POI Prediction (NPP): This task converts a POI-based trajectory into a sequence of representation vectors using an encoder. The downstream model, often a deep sequential model like LSTM, predicts the next POI in the trajectory using the representation of the first half. The output is a multi-class classification, with each class representing a possible POI.
  • Trajectory User Link (TUL): This task predicts the user who generated a trajectory based on the trajectory’s representation sequence. The downstream model is a deep sequential model, with each output class corresponding to a user ID.
The fine-tuning loss for all three tasks is the cross-entropy loss.
@@
  • Average Speed Inference (ASI). This task leverages the representation vector to infer the average speed of a road segment. The average speed is derived from trajectory data and must be excluded from the input features of the representation encoders to prevent data leakage. Inferring average speed is a representative task for estimating unknown features of a road segment. The downstream model for this task is a linear regression function, and the fine-tuning loss function is Mean Square Error (MSE).
  • Travel Time Estimation (TTE). This task estimates the travel time from the origin to the destination along a trajectory composed of road segments. The downstream model is a deep sequential model. Its input is the representation sequence of the segment-based trajectory, and the output is the regressed travel time. The fine-tuning loss function is MSE.
  • Similarity Trajectory Search (STS). This task aims to find the most similar trajectory from a large database given a trajectory query. The downstream model first employs a deep sequential model to convert the representation sequences of all trajectories in the database into trajectory representation vectors. Then, it uses cosine similarity between these representations to identify the most similar trajectory to the query. The fine-tuning loss function is a contrastive learning loss, such as InfoNCE, which treats the ground-truth trajectory as the positive sample and others as negative samples. The ground-truth trajectory is typically generated by replacing a randomly selected sub-trajectory with another that shares the same start and endpoints (Jiang et al., 2023; Mao et al., 2022; Chang et al., 2023).
@@
  • Land Parcel Classification (LPC): This task predicts specific labels for a land parcel, such as land usage or functional category, based on its representation vector. The downstream function is typically a MLP with a SoftMax classifier. The fine-tuning loss is cross-entropy.
  • Flow Inference (FI): This task predicts the incoming and outgoing population of a land parcel over a given period (e.g., one month) using its representation vector. The downstream function is usually a linear regression, with MSE as the loss.
  • Mobility Inference (MI): This task predicts the OD flow between pairs of land parcels. The input to the downstream function consists of the representations of two parcels, and the output is the predicted OD flow volume. The downstream function can be a neural network, with MSE as the fine-tuning loss.
@@ 3.1. Data Module and Atomic Files
In the Data Module, map entities and auxiliary data are represented in a unified file format, called Atomic Files (Wang et al., 2021). This format consists of three file types: .geo -– represents map entities, .traj -– stores trajectory data, .rel – captures relation networks.
.geo Files. A .geo file stores a list of map entities, such as POIs, road segments, and land parcels. Each line in the file represents an entity with four fields: geo_uid, geo_type, geo_location, and geo_features.
  • geo_uid. This field records the unique identifier of a map entity.
  • geo_type. This field indicates the entity type, which can be “point,” “polyline,” or “polygon,” corresponding to POIs, road segments, and land parcels, respectively.
  • geo_location. This field stores the geographical coordinates of the entity: a single longitude-latitude pair for points, a sequence of coordinates for polylines, and a closed sequence for polygons.
  • geo_features (optional). This field contains the entity’s features, with its format and length varying based on the data type and number of features.
.traj Files. A .traj file stores a set of map entity-based trajectories. Each line corresponds to a trajectory sample and contains four fields: traj_uid, sample_id, entity_id, and time_stamp.
  • traj_uid. Specifies the unique ID of the trajectory to which the sample belongs.
  • sample_order. Records the sequential index of the sample within the trajectory.
  • entity_id. Identifies the map entity corresponding to the sample, with its value matching the geo_uid in the .geo file.
  • time (optional). Stores the timestamp of the sample in Unix timestamp format.
.rel Files. The .rel file uses a sparse matrix format to represent the adjacency matrix of a relational network. Each line records a non-zero relationship between two map entities, containing the following fields: rel_uid, orig_geo_id, dest_geo_id, weight, feature, and timestamp.
  • rel_uid. Records the unique ID for the relationship between two map entities.
  • orig_geo_id and dest_geo_id. Specify the unique IDs of the origin and destination map entities involved in the relation, matching the geo_uid values in the .geo file.
  • weight. Stores the weight of the edge corresponding to the map entity relation.
  • feature (optional). Contains additional features for the edge. In some models, relation networks are represented as heterogeneous graphs, and this field can store the type of edges.
  • time (optional). Stores the timestamp for the relation, which is essential for dynamic graphs where edge weights and features may change over time.
Using the three types of atomic files, we can represent map entities and auxiliary data in a unified format. Original data often comes in various spatiotemporal formats, such as GeoJSON (Butler et al., 2016), Shapefile (Environmental Systems Research Institute, 1998), and NPZ (Harris et al., 2020), requiring multiple data-loading implementations and hindering reusability. Our VecCity library solves this issue by adopting unified atomic files, simplifying data processing.
@@
3.2. Upstream Module and Downstream Module
  • encode(): This interface implements the encoder for an MapRL model. It takes data extracted from atomic files as input and outputs representation vectors for map entities.
  • pretraining_loss(): This interface manages the implementation of pre-training tasks. It prepares samples through methods such as data masking, generating positive/negative samples, or data augmentation, and calculates the corresponding loss values.
In the VecCity framework, users only need to implement the interface functions in the upstream module to execute the pre-training phase of an MapRL model. Using these two interfaces, we have reproduced 21 mainstream MapRL models in VecCity, with 7 models each for POI, road segment, and land parcel representation learning. The names of the reproduced models are marked with “*” in Tab. 1.
  • downstream_model(): This interface implements the downstream models used during the fine-tuning step. It typically utilizes a simple MLP or LSTM model. For classification tasks, the output layer is a SoftMax classifier, while for continuous prediction tasks, it is a linear regression layer.
  • finetuning_loss(): This interface defines the loss function for fine-tuning. Cross-entropy is used for classification tasks, MSE for regression tasks, and InfoNCE for contrastive learning tasks.
Table 3. Performance comparison of POI MapRL methods on various datasets. The bold/underlined font means the best/the second-best result. "↑" means bigger is better and "↓" means smaller is better.
Dataset New York Tokyo Chicago Singapore San Francisco Overall
Task POIC NPP POIC NPP POIC NPP POIC NPP POIC NPP Pre Ave Rank Pre Rank
Metric ACC@1↑ F1↑ ACC@1↑ ACC@5↑ ACC@1↑ F1↑ ACC@1↑ ACC@5↑ ACC@1↑ F1↑ ACC@1↑ ACC@5↑ ACC@1↑ F1↑ ACC@1↑ ACC@5↑ ACC@1↑ F1↑ ACC@1↑ ACC@5↑
SkipG 0.071 0.012 0.126 0.309 0.334 0.042 0.164 0.333 0.077 0.007 0.050 0.161 0.027 0.003 0.030 0.074 0.011 0.003 0.014 0.053 5.3 7
Tale 0.057 0.009 0.152 0.341 0.212 0.037 0.145 0.297 0.096 0.009 0.094 0.230 0.029 0.005 0.043 0.101 0.009 0.001 0.007 0.046 4.8 6
Teaser 0.079 0.014 0.131 0.329 0.354 0.048 0.137 0.241 0.077 0.005 0.029 0.137 0.065 0.018 0.033 0.078 0.037 0.012 0.014 0.058 4.4 5
Hier 0.052 0.008 0.121 0.285 0.278 0.032 0.162 0.335 0.086 0.011 0.036 0.161 0.041 0.006 0.079 0.163 0.039 0.007 0.012 0.108 4.2 4
P2Vec 0.074 0.013 0.116 0.154 0.313 0.037 0.154 0.319 0.111 0.012 0.143 0.393 0.068 0.014 0.075 0.168 0.033 0.011 0.019 0.046 3.7 3
CACSR 0.028 0.004 0.135 0.331 0.232 0.010 0.159 0.325 0.104 0.017 0.146 0.394 0.030 0.005 0.103 0.126 0.018 0.004 0.077 0.158 3.5 2
CTLE 0.048 0.006 0.154 0.336 0.235 0.014 0.168 0.338 0.119 0.019 0.187 0.398 0.083 0.008 0.107 0.224 0.036 0.006 0.060 0.123 2.0 1
Task TUL Efficiency TUL Efficiency TUL Efficiency TUL Efficiency TUL Efficiency Eff Ave Rank Eff Rank
Metric ACC@1↑ F1↑ Param↓ Speed↓ ACC@1↑ F1↑ Param↓ Speed↓ ACC@1↑ F1↑ Param↓ Speed↓ ACC@1↑ F1↑ Param↓ Speed↓ ACC@1↑ F1↑ Param↓ Speed↓
SkipG 0.625 0.317 2.632 2 0.452 0.243 2.013 3 0.314 0.085 0.937 1 0.158 0.047 2.513 3 0.123 0.138 0.414 1 1.0 1
Tale 0.699 0.429 8.611 5 0.395 0.219 8.111 12 0.609 0.360 2.910 2 0.300 0.164 9.919 13 0.329 0.151 1.107 1 4.4 6
Teaser 0.608 0.307 3.012 3 0.464 0.255 2.318 5 0.312 0.193 1.064 1 0.173 0.152 2.986 5 0.556 0.184 0.464 1 1.8 2
Hier 0.635 0.347 8.253 5 0.493 0.251 7.631 8 0.521 0.289 6.550 2 0.136 0.142 8.133 9 0.557 0.242 6.025 1 4.2 4
P2Vec 0.629 0.323 7.142 7 0.451 0.243 2.735 12 0.647 0.413 1.410 2 0.386 0.169 3.346 11 0.517 0.256 1.496 1 3.6 3
CACSR 0.645 0.431 9.921 43 0.502 0.249 8.301 114 0.659 0.462 5.439 21 0.361 0.156 9.715 68 0.611 0.295 4.059 7 6.7 7
CTLE 0.698 0.444 3.436 8 0.507 0.294 2.814 13 0.772 0.628 1.734 4 0.485 0.243 3.316 12 0.628 0.323 1.209 1 4.2 4

### v2

@@ 2.2.2. Graph-based Models
Method Taxnomy Encoders Pretraining Tasks Map Entity Features Aux. Data
Category Token Graph Seq TokRI TRCL AToCL NFI GAu NCL AGCL TrajP MTR ATrCL POI RS LP Sem Spa Temp Traj Rel
Point of Interest CWAP (Liu et al., 2016) Token-based CIT
CAPE (Chang et al., 2018) Token-based CIT
MT-POI (Yao et al., 2023) Token-based GR
Place2V (Yan et al., 2017) Token-based GR
Semant (Huang et al., 2022) Token-based GR
CatEM (Bing et al., 2023) Token-based CIT
SkipG* (Mikolov et al., 2013) Token-based CIT
Tale* (Wan et al., 2021) Token-based CIT
Teaser* (Zhao et al., 2017) Token-based CIT GR
Hier* (Shimizu et al., 2020) Seq.-based CIT GR
P2Vec* (Feng et al., 2017) Token-based CIT GR
CACSR* (Gong et al., 2023) Seq.-based CIT
CTLE* (Lin et al., 2021) Seq.-based CIT
Road Segment DyToast (Chen et al., 2024b) Graph-based CDT GR
R2Vec* (Jepsen et al., 2018) Token-based GR
TrajRNE (Schestakov et al., 2023) Graph-based CDT GR/SR
JGRM (Ma et al., 2024) Seq.-based CDT GR
RN2Vec* (Wang et al., 2020) Token-based GR
HRNR* (Wu et al., 2020) Graph-based GR
SARN* (Chang et al., 2023) Graph-based GR
Toast* (Chen et al., 2021) Seq.-based CDT GR
HRoad* (Zhang and Long, 2023) Graph-based GR
START* (Jiang et al., 2023) Seq.-based CDT GR
JCLRNT* (Mao et al., 2022) Seq.-based CDT GR
Land Parcel CGAL (Zhang et al., 2019a) Graph-based GR/SR
DLCL (Du et al., 2019) Graph-based GR/SR
HUGAT (Kim and Yoon, 2022) Graph-based GR/SR
Re2Vec (Luo et al., 2022) Graph-based GR/SR
ReCP (Li et al., 2024a) Token-based SR
HGI (Huang et al., 2023) Graph-based GR
ReDCL (Li et al., 2023) Token-based GR
HAFus (Sun et al., 2024) Graph-based GR/SR
ZEMob* (Yao et al., 2018) Token-based SR
GMEL* (Liu et al., 2020) Graph-based GR/SR
MGFN* (Wu et al., 2022) Graph-based SR
HDGE* (Wang and Li, 2017) Token-based GR/SR
MVURE* (Zhang et al., 2020) Graph-based SR
ReMVC* (Zhang et al., 2022) Token-based SR
HREP* (Zhou et al., 2023) Graph-based GR/SR
@@ 2.2.3. Sequence-based Models
Encoders. Sequence-based encoders are designed to capture temporal dependencies among map entities within trajectory-type auxiliary data. They employ sequential deep learning models, such as LSTMs (Sak et al., 2014), GRUs (Cho, 2014), or Transformers (Vaswani et al., 2017), to encode the embeddings of map entities along a trajectory into a sequence of representation vectors. These input embeddings are typically derived from token-based or graph-based encoders. Specifically, for a trajectory $tr$ consisting of $K$ entities, let their embedding sequence be $\bm{R}_{tr}=(\bm{r}_{tr_{1}},\ldots,\bm{r}_{tr_{k}},\ldots,\bm{r}_{tr_{K}})$. The sequence-based model then processes this input to generate corresponding representation vectors as $\bm{S}_{tr}=(\bm{s}_{tr_{1}},\ldots,\bm{s}_{tr_{k}},\ldots,\bm{s}_{tr_{K}})=\mathrm{SEQ}(\bm{R}_{tr}),$ where $\mathrm{SEQ}(\cdot)$ denotes the sequence-based encoder, and $\bm{s}_{tr_{k}}$ represents the embedding of the $k$-th entity along the trajectory $tr$.
In trajectory data, each location sample is associated with a timestamp. Integrating temporal information from timestamps into entity representations has become a key focus of sequence-based encoders. Early methods divide time into discrete slots, such as the hour of the day or the day of the week, and encoding these slots as embeddings using token-based encoders (Mao et al., 2022; Feng et al., 2017; Zhao et al., 2017). More advanced models embed timestamps within the components of sequential models. For example, JCLRNT (Mao et al., 2022), Toast (Chen et al., 2021), and CTLE (Lin et al., 2021) incorporate timestamps into the positional encoding of Transformers. Similarly, START (Jiang et al., 2023) integrates travel time into the attention mechanism, enhancing the model’s ability to capture temporal dependencies.
@@
$\bullet$ Trajectory Prediction (TrajP). The goal of this task is to predict the second half of a trajectory using the first half. Given a trajectory $tr$, this task is formally defined as $\hat{tr}_{[k+1:K]}=Pre(\bm{S}_{[1:k]}),~{}\bm{S}_{[1:k]}=Enc(tr_{[1:k]})$, where $Enc(\cdot)$ is the sequence-based encoder, and $Pre(\cdot)$ is the prediction function. Here, $tr_{[1:k]}$ represents the first $k$ samples of the trajectory, and $\bm{S}_{[1:k]}$ is their corresponding sequence of representation vectors. The objective is to minimize the distance between $\hat{tr}_{[k+1:K]}$ and the ground truth.
[…]
$\bullet$ Augmented Trajectory Contrastive Learning (ATrCL). This task generates multiple versions of a trajectory using data augmentation techniques (Fang et al., 2020; Zhao et al., 2019). Sequence-based encoders then encode both the original and augmented trajectories, and a contrastive learning loss is applied to enhance the similarity between representations of the same trajectory while distinguishing them from others.
In TrajP tasks, the prediction labels can be various. For example, Hier (Shimizu et al., 2020) uses entity IDs as prediction labels, while CTLE (Lin et al., 2021) employs visiting time as labels to incorporate temporal information. In MTR tasks, different strategies exist for selecting masked positions. A basic approach is to mask positions within trajectories (Lin et al., 2021) randomly, but this overlooks the sequential dependencies between samples. To address this, START (Jiang et al., 2023) and Toast (Chen et al., 2021) mask continuous subsequences within a trajectory. In ATrCL tasks, models adopt diverse augmentation strategies. JCLRNT (Mao et al., 2022) generates augmented trajectories by randomly deleting or replacing trajectory points. Considering contextual dependencies, START (Jiang et al., 2023) replaces continuous sub-sequences, while CACSR (Gong et al., 2023) injects white noise into representations to enhance robustness.
@@ 2.2.4. Module Integration of Encoders and Pre-training Tasks
The encoders and pre-training tasks introduced in the previous sections serve as the foundational components of MapRL models. A complete MapRL model can integrate multiple types of encoders and pre-training tasks, leveraging their complementary strengths to enhance representation learning. For the encoder modules, existing models typically follow a “token $\rightarrow$ graph $\rightarrow$ sequence” modeling pipeline. This pipeline first extracts basic features using token-based encoders, then captures relational structures through graph-based encoders, and finally models temporal dependencies with sequence-based encoders. At each stage of this pipeline, a MapRL model may employ multiple encoders simultaneously for different pre-training tasks. Table 1 provides an overview of how mainstream models combine encoders and pre-training tasks.
For the three-stage encoder pipeline, MapRL models follow two pre-training paradigms. $i$) Sequential Training: Encoders are trained in stages—first token-based, then graph-based, and finally sequence-based—where each stage refines the output of the previous one. $ii$) Joint Training: All encoders are optimized simultaneously using multi-task learning, enabling the model to capture entity features, structural relationships, and temporal patterns in an integrated manner. These paradigms offer flexibility in model impanelments, depending on the specific requirements of the MapRL task.
@@ 2.3. Downstream Tasks
Tasks for POIs. Common downstream tasks for POI entities can be classified into three categories, all of which utilize cross-entropy loss as the objective function:
$\bullet$ POI Classification (POIC): This task aims to classify a POI into predefined categories based on its representation. A typical downstream model consists of an MLP followed by a SoftMax classifier.
$\bullet$ Next POI Prediction (NPP): This task converts a POI-based trajectory into a sequence of representation vectors using an encoder. The downstream model, often a deep sequential model like LSTM, predicts the next POI in the trajectory using the representation of the first half. The output is a multi-class classification, with each class representing a possible POI.
$\bullet$ Trajectory User Link (TUL): This task predicts the user who generated a trajectory based on the trajectory’s representation sequence. The downstream model is a deep sequential model, with each output class corresponding to a user ID.
@@
$\bullet$ Average Speed Inference (ASI). This task leverages the representation vector to infer the average speed of a road segment. The average speed is derived from trajectory data and must be excluded from the input features of the representation encoders to prevent data leakage. Inferring average speed is a representative task for estimating unknown features of a road segment. The downstream model for this task is a linear regression function, and the fine-tuning loss function is Mean Square Error (MSE).
$\bullet$ Travel Time Estimation (TTE). This task estimates the travel time from the origin to the destination along a trajectory composed of road segments. The downstream model is a deep sequential model. Its input is the representation sequence of the segment-based trajectory, and the output is the regressed travel time. The fine-tuning loss function is MSE.
$\bullet$ Similarity Trajectory Search (STS). This task retrieves the most similar trajectory from a database given a query. A deep sequential model encodes all trajectories into representation vectors, and cosine similarity identifies the closest match. Fine-tuning uses a contrastive loss (e.g., InfoNCE), treating the ground-truth trajectory—formed by replacing a sub-trajectory with one sharing the same endpoints (Jiang et al., 2023; Mao et al., 2022; Chang et al., 2023)—as the positive sample.
@@
$\bullet$ Land Parcel Classification (LPC): This task predicts specific labels for a land parcel, such as land usage or functional category, based on its representation vector. The downstream model is an MLP with a SoftMax classifier, optimized with cross-entropy loss.
$\bullet$ Flow Inference (FI): This task predicts the incoming and outgoing population of a land parcel over a given period (e.g., one month) using its representation vector. The downstream function is usually a linear regression, with MSE as the loss.
$\bullet$ Mobility Inference (MI): This task predicts the OD flow between pairs of land parcels. The input to the downstream function consists of the representations of two parcels, and the output is the predicted OD flow volume. The downstream function can be a neural network, with MSE as the fine-tuning loss.
@@ 3.1. Data Module and Atomic Files
In the Data Module, map entities and auxiliary data are represented in a unified file format, called Atomic Files (Wang et al., 2021). This format consists of three file types: .geo – represents map entities, .traj – stores trajectory data, .rel – captures relation networks. Using the three types of atomic files, we can represent map entities and auxiliary data in a unified format. Original data often comes in various spatiotemporal formats, such as GeoJSON (Butler et al., 2016), Shapefile (Environmental Systems Research Institute, 1998), and NPZ (Harris et al., 2020), requiring multiple data-loading implementations and hindering reusability. Our VecCity library solves this issue by adopting unified atomic files, simplifying data processing.
Figure 2. The overall pipeline for VecCity.
@@
3.2. Upstream Module and Downstream Module
$\bullet$ encode(): This interface implements the encoder for an MapRL model. It takes data extracted from atomic files as input and outputs representation vectors for map entities.
$\bullet$ pretraining_loss(): This interface manages the implementation of pre-training tasks. It prepares samples through methods such as data masking, generating positive/negative samples, or data augmentation, and calculates the corresponding loss values.
In the VecCity framework, users can execute the pre-training phase of an MapRL model by implementing the interface functions in the upstream module. Leveraging these interfaces, we have reproduced 21 mainstream MapRL models in VecCity, including 7 models each for POI, road segment, and land parcel representation learning. The reproduced models are marked with “*” in Tab. 1.
$\bullet$ downstream_model(): This interface implements the downstream models used during the fine-tuning step. It typically utilizes a simple MLP or LSTM model. For classification tasks, the output layer is a SoftMax classifier, while for continuous prediction tasks, it is a linear regression layer.
$\bullet$ finetuning_loss(): This interface defines the loss function for fine-tuning. Cross-entropy is used for classification tasks, MSE for regression tasks, and InfoNCE for contrastive learning tasks.